------------------------------------------------------------ Date: Fri Nov 2 04:38:09 2001, Shift: 00:00-08:00 Supervisor: DO, with: Log book#10 pages 85 - 86 Shift Summary: ------------------------------------------------------------ Early report for this shift! 1.Machine status: bad news for Brahms. The last couple of stores, although good for the other experiments were totally useless for us, we never had enough coincidences to take useful data (rate between 0 and 10 most of the time) I called MCR many times to ask about it, they tried to investigate the problem, improve the steering for us, etc but finally, they gave up, they explicitely told me that the next stores won't be useful for us and that I should rather go home (on the other hand, Phenix has now a huge rate, I saw 5400!! 15 mn ago). 2.Achievements: Small data run of 20 mn when the ZDC coincindence was about 200 Hz after a steering that MCR could not keep. We got only 6.4k trigger 6 in run 5627. Otherwise, DVM run (5627), TPC pedestals (5623), Pedestal run (5628) 3.Nothing to declare. Note: I'll leave the magnets off. MCR is aware. ------------------------------------------------------------
This archive was generated by hypermail 2b30 : Fri Nov 02 2001 - 04:38:45 EST